-2s + 6 = -5s + 12
+2s +2s
6 = -3s + 12
- 12 - 12
- 6 = -3s
6 = 3s
6/3 = 3s/3
2 = s

Answer by Diana541(24)

-2s + 6 = -5s + 12
+2s +2s
6 = -3s + 12
- 12 - 12
- 6 = -3s
6 = 3s
6/3 = 3s/3
2 = s